C8H10N4O2 percent composition



1
Expert's answer
2021-02-24T03:50:01-0500

w=(n*Ar/Mr)*100%

Mr=8*12+10*1+4*14+2*16=96+10+56+32=194

w(C)=(96/194)*100%=49.48%

w(H)=(10/194)*100%=5.15%

w(N)=(56/194)*100%=28.87%

w(O)=(32/194)*100%=16.49%